# MCPDepth: Omnidirectional Depth Estimation via Stereo Matching from Multi-Cylindrical Panoramas

### v1

@@ Abstract
We introduce Multi-Cylindrical Panoramic Depth Estimation (MCPDepth), a two-stage framework for omnidirectional depth estimation via stereo matching between multiple cylindrical panoramas. MCPDepth uses cylindrical panoramas for initial stereo matching and then fuses the resulting depth maps across views. A circular attention module is employed to overcome the distortion along the vertical axis. MCPDepth exclusively utilizes standard network components, simplifying deployment to embedded devices and outperforming previous methods that require custom kernels. We theoretically and experimentally compare spherical and cylindrical projections for stereo matching, highlighting the advantages of the cylindrical projection. MCPDepth achieves state-of-the-art performance with an 18.8% reduction in mean absolute error (MAE) for depth on the outdoor synthetic dataset Deep360 and a 19.9% reduction on the indoor real-scene dataset 3D60.
@@ 1 Introduction
Figure 1: Comparison between MODE and our method. The left part of the figure compares the performance of MODE and our method on the stereo matching task. The first and second columns display the results from MODE and our method, respectively. (a) and (b) show spheres and spherical stereo panoramas. (c) presents disparity ground truth and (d) the estimated disparity map from MODE. (e) and (f) depict cylinders and cylindrical stereo panoramas. (g) shows the estimated disparity map from our method, and (h) displays the converted disparity map in Cassini projection from (g), we evaluate the corresponding center area in (d). On the right part, (i), (j), and (k) represent the ground truth of the depth map, the depth map from MODE, and the depth map from our method, respectively. Our method outperforms MODE in both stereo matching and depth estimation tasks.
Depth estimation is a fundamental challenge in geometric computer vision. Omnidirectional depth estimation has recently gained tremendous progress with the advent of convolutional neural networks and has been widely applied in robot perception. Various methods have been proposed for both monocular [44, 15] and binocular [46, 19, 20] settings. Some utilize regular convolution for spherical projection [44, 15, 46], while others do not simplify the spherical epipolar constraint [19]. Additionally, some methods employ customized convolutions, such as deformable convolution [41], EquiConvs [9], and spherical convolution [20], which pose deployment challenges [53, 38] on robot platforms. Furthermore, depth estimation from one or two views is unreliable and error-prone.
Several works have explored the multi-view setting to overcome these challenges, such as SweepNet [51] and OmniMVS [50], utilizing four fish-eye cameras. However, these methods struggle with effective feature extraction due to distortion and large field-of-view (FoV) as they rely on regular convolutions. In addition, the blind spots in fish-eye cameras contribute to discontinuities in the spherical cost volume.
MODE [20] introduces a two-stage framework of omnidirectional depth estimation from multi-view spherical panoramas and proposes the Deep360 dataset. MODE adopts spherical convolutions [7] to address the distortion problem in spherical panoramas, encountering deployment challenges [38]. MODE also employs Cassini projection [48] for stereo matching to simplify the epipolar constraint. Cassini projection, the transverse aspect of the equirectangular projection (ERP), is widely used in map projection. However, compared to cylindrical projection [1], Cassini projection has greater distortion, which increases the closer it gets to the poles.
We adopt the two-stage structure from MODE [20] but use multi-view panoramas in cylindrical projection. This projection reduces distortion, avoids customized convolutions, and maintains a simple epipolar constraint. Fig. 1 illustrates the differences between Cassini and cylindrical projections for stereo matching and their respective performances. We also propose a circular attention module to better capture panorama features with a $360^{\circ}$ vertical FoV and overcome vertical axis distortion.
Our contributions can be summarized as follows:
@@
  • We present the first omnidirectional depth estimation framework via stereo matching between multiple cylindrical panoramas.
  • We theoretically and experimentally compare spherical and cylindrical projections for stereo matching, summarizing the advantages of cylindrical projection.
  • We propose a circular attention module to address vertical axis distortion in cylindrical panoramas and enhance the receptive fields of convolutional neural networks (CNNs).
  • Our approach achieves state-of-the-art performance on the outdoor synthetic dataset Deep360 and the indoor real-scene dataset 3D60.
@@ 2.1 Deep Learning-based Stereo Matching
Early methods used deep neural networks to compute matching costs, like MCCNN [58], which trains a CNN for initial patch matching costs. Recently, end-to-end neural networks have dominated the stereo matching methods. Works such as [28, 30, 23, 11, 55, 22, 40] only use 2D convolutions. Mayer et al. [28] propose the first end-to-end disparity estimation network, DispNet, and its correlation version, DispNetC. Pang et al. [30] introduce a two-stage framework named CRL with multi-scale residual learning. GwcNet [11] proposes the group-wise correlation volume to improve the expressiveness of the cost volume and performance in ambiguous regions.
AANet [55] adopts a novel aggregation algorithm using sparse points and multi-scale interaction. Another series of works [16, 3] use 3D convolutions, which have great potential in regularizing or filtering the cost volume. GCNet [16] firstly uses 3D encoder-decoder architecture to regularize a 4D concatenation volume. PSMNet [3] proposes a stacked hourglass 3D CNN in conjunction with intermediate supervision to regularize the concatenation volume.
Recently, iterative methods [25, 43, 18, 54] have shown impressive results. RAFTStereo [25] proposes to recurrently update the disparity field using local cost values retrieved from the all-pairs correlations. IGEV-Stereo [54] introduces a geometry encoding volume to encode non-local geometry and context information.
Significant advancements have also been made in multi-view stereo (MVS) [57, 56, 4, 10]. Note that these methods are designed for perspective cameras with normal FoV.
In this work, we base our experiments on the cost volume-based method PSMNet and the iterative method IGEV-Stereo.
@@ 2.2 Omnidirectional Depth Estimation
Omnidirectional depth estimation has developed tremendously with neural networks. Zioulis et al. present a learning-based monocular depth estimation method in  [61], trained directly on omnidirectional content in the ERP domain and later propose CoordNet [62] with spherical disparity model. BiFuse [44] uses both equirectangular and cubemap projections for depth estimation. A more effective fusion framework for ERP and cubemap projection is proposed in Unifuse [15]. Cheng et al. [5] introduce a depth sensing system by combining an OmniCamera with a regular depth sensor. 360SD-Net [46] is the first end-to-end trainable network for stereo depth estimation using spherical panoramas. CSDNet [19] focuses on the left-right stereo and uses Mesh CNNs [14] to overcome the spherical distortion. SweepNet [51] and OmniMVS [50] use multi-view fish-eye images for omnidirectional depth maps. However, most of these methods are based on spherical projection and extract spherical features with regular convolutions and none of them discuss the properties of cylindrical projection.
Cheng et al. [5] propose a spherical feature transform layer to reduce the difficulty of feature learning. MODE [20] adopts spherical convolution from Spherenet [7], but the customized CUDA implementation poses deployment challenges on robotic computing platforms [38].
Jun et al. [37] employed cylindrical panoramas for stereo matching, but without CNNs or analysis of cylindrical projection properties, and the panoramas were stitched from 12 perspective images. MCPDepth is the first to explore the advantages of cylindrical projection over spherical projection for stereo matching and to use cylindrical panoramas for omnidirectional depth estimation.
@@ 2.3 Self-Attention Module
Attention mechanisms are first introduced by [2] for the encoder-decoder in a neural sequence-to-sequence model. It is designed to capture the correspondence of tokens between two sequences. Self-attention, designed for single contexts, can encode long-range interactions. Recently, self-attention has been widely applied in computer vision and achieves state-of-the-art performance [42, 31, 47, 59, 13, 29, 32, 39, 6].
In stereo matching, CREStereo [18] first adopts the self-attention module from LoFTR [39]. Zhao et al. [60] propose a multi-stage and multi-scale channel-attention transformer to preserve high-frequency information. GOAT [26] uses self-cross attention to capture more representative and distinguishable features. However, these methods are not designed for stereo matching of panoramas with a $360^{\circ}$ FoV.
Global self-attention in image processing is computationally expensive due to the need to calculate the relationship between every pixel and every other pixel. This limits its practical usage across all layers in a full-attention model. It is shown in [35, 12] that self-attention layers alone could be stacked to form a fully attentional model by restricting the receptive field of self-attention to a local square region. Inspired by Axial-DeepLab [45], we propose circular attention to capture the $360^{\circ}$ features. We apply circular attention only along the vertical axis of feature maps, which significantly reduces the training resources.
@@ 3 Method
3.1 Framework
Given $m$ $360^{\circ}$ cameras, where $m\geq 3$, we have a set of $n=\mathrm{C}_{2}^{m}=\frac{m!}{2!(m-2)!}$ pairs of rectified panoramas $\{(I^{i}_{L},I^{i}_{R})\}^{n}_{i=1}$ with their extrinsic parameters. Our goal is to estimate the omnidirectional depth map $d$ for the left panorama in the first pair $I^{1}_{L}$. The MCPDepth framework, shown in Fig. 2, includes two stages. In the stereo matching stage, $n$ pairs of rectified cylindrical panoramas (Fig. 2 (a)) are fed into the stereo matching network. The number of pairs ($n$) and cameras ($m$) varies on different datasets: $n=6,m=4$ for Deep360, and $n=3,m=3$ for 3D60. The resulting disparity and confidence maps (Fig. 2 (b)) are reprojected into the Cassini domain with a $180^{\circ}$ horizontal FoV. The disparity maps are then converted to depth maps. The depth and confidence maps are aligned with the view of $I^{1}_{L}$ using extrinsic parameters as shown in Fig. 2 (c). Black areas indicate invisible and occluded regions.
We generally follow MODE’s depth fusion stage structure. Specifically, multi-view depth maps, along with their corresponding confidence maps and reference panoramas, are inputted into two separate 2D encoder blocks. The fused depth map is then processed through a single decoder block, incorporating skip connections between the encoder and decoder blocks at each scale. The final depth map is in the Cassini domain but can be converted to the ERP domain. More details are available in the supplementary material.
3.2 Stereo Matching
We use the circular attention module between feature extraction and cost volume with a structure similar to PSMNet [3]. The circular attention module augments the extracted features to capture features from a $360^{\circ}$ FoV and overcome vertical axis distortion. These augmented features are then shifted and concatenated to build the cost volume. The disparity map is regressed through the 3D stacked hourglass network. During training, we use the L1 loss to train the network. The confidence maps are used to measure the reliability of the disparity estimation and are widely used in stereo matching tasks [34]. The confidence map is obtained during inference. Specifically, considering the disparity is obtained through a probability-weighted sum over all disparity hypotheses, we compute the corresponding confidence value by taking a probability sum over the three nearest disparity hypotheses.
Cylindrical Projection Fig. 3 highlights the differences between spherical projection and cylindrical projection for stereo matching. Cylindrical projection maintains the linear epipolar constraint.
@@ 3.2 Stereo Matching
Figure 3: (a) and (b) compare the spherical and cylindrical projections for stereo matching and their respective epipolar geometries. (c) and (d) represent the schematic drawing of the epipolar plane under spherical and cylindrical projections. (e) and (f) show the different sampling patterns on the sphere and the cylinder when regular convolutions are applied to spherical and cylindrical panoramas.
@@
Eq. 1 illustrates the conversion between spherical and cylindrical coordinate systems and the Cartesian coordinate system. In spherical coordinates, $\rho$ is the Euclidean distance from the origin $O$ to point $P$; $\phi$ is the angle between line $OP$ and the plane $yOz$; $\theta$ is the angle between line $OP^{\prime}$ and the z-axis, where $P^{\prime}$ is the projection of $P$ on the plane of $yOz$. In cylindrical coordinates, $\rho$ is the Euclidean distance from the x-axis to the point $P$; $\theta$ is the angle between line $OP^{\prime}$ and the z-axis, where $P^{\prime}$ is the projection of $P$ on the plane of $yOz$.
[…]
@@
[…]
In distortion-free perspective images, an object’s actual length and its pixel length along the horizontal and vertical axes is given by $\Delta u=\frac{f_{x}}{z}\Delta x,\Delta v=\frac{f_{y}}{z}\Delta y$, where $f_{x},f_{y}$ are the focal lengths along the $x$ and $y$ axes, the $z$ is the distance along the z-axis. Eq. 3 shows these relationships for both spherical and cylindrical projections.
[…]
where all focal lengths are $f=R=H/2\pi$ in this work. The approximation of $\Delta X\approx\rho\cos{\theta}\Delta\phi$, $\Delta Y\approx\rho\Delta\theta$ is under the condition that the object is not too large or far from the camera [33]. This approximation means objects in cylindrical projection appear similar regardless of their location. This shift-invariant property facilitates efficient learning by CNNs. In contrast, objects in spherical projection vary with their $\theta$ axis position, which limits the effectiveness of regular convolutions, as shown in Fig. 3 (e).
@@
where $B$ denotes the baseline between left and right origins. As shown in Fig. 3 (d), the cylindrical projection maintains the same disparity-depth relationship as perspective images:
$\rho_{l}=\frac{B\cdot R}{\left|x_{l}-x_{r}\right|}$ (5)
@@
The benefits of cylindrical projection for stereo matching of panoramas can be summarized as (1) Compatibility with perspective images: Disparity in cylindrical panoramas is defined as $d=\left|x_{l}-x_{r}\right|$, matching the disparity in perspective images. This allows stereo matching networks designed for perspective images to be directly applicable. (2) Reduced distortion: Spherical projection introduces distortion along both the vertical and horizontal axes, with distortion increasing near the poles. In contrast, cylindrical projection exhibits distortion solely along the vertical axis. This better shift-invariant property shown in Fig. 3 (f) enables CNNs to learn features more efficiently. (3) Simplified Deployment: Spherical panoramas require customized convolutions [41, 7, 9] to extract features. For example, spherical convolutions in MODE can’t be exported to widely used ONNX [49] format for deployment, they either need CUDA Plugins for the TensorRT engine on NVIDIA platforms or customized implementation on other embedded devices. Cylindrical panoramas only use regular convolutions, making MCPDepth deployment-friendly.

### v2

# MCPDepth: Practical Omnidirectional Depth Estimation from Multiple Cylindrical Panoramas via Stereo Matching

@@ Abstract
Omnidirectional depth estimation presents a significant challenge due to the inherent distortions in panoramic images. Despite notable advancements, the impact of projection methods remains underexplored. We introduce Multi-Cylindrical Panoramic Depth Estimation (MCPDepth), a novel two-stage framework designed to enhance omnidirectional depth estimation through stereo matching across multiple cylindrical panoramas. MCPDepth initially performs stereo matching using cylindrical panoramas, followed by a robust fusion of the resulting depth maps from different views. Unlike existing methods that rely on customized kernels to address distortions, MCPDepth utilizes standard network components, facilitating seamless deployment on embedded devices while delivering exceptional performance. To effectively address vertical distortions in cylindrical panoramas, MCPDepth incorporates a circular attention module, significantly expanding the receptive field beyond traditional convolutions. We provide a comprehensive theoretical and experimental analysis of common panoramic projections—spherical, cylindrical, and cubic—demonstrating the superior efficacy of cylindrical projection. Our method improves the mean absolute error (MAE) by 18.8% on the outdoor dataset Deep360 and by 19.9% on the real dataset 3D60. This work offers practical insights for other tasks and real-world applications, establishing a new paradigm in omnidirectional depth estimation. The code is available at https://github.com/Qjizhi/MCPDepth.
@@ 1 Introduction
Figure 1: Comparison of stereo images among common panoramic projections.
Depth estimation is a pivotal challenge in geometric computer vision, playing a critical role in 3D scene understanding and robotic perception. Despite substantial advancements achieved through convolutional neural networks (CNNs) in processing perspective images, the task of estimating omnidirectional depth remains particularly challenging due to the severe geometric distortions inherent in panoramic representations. Recent research has investigated both monocular [45, 14] and stereo [47, 18, 19] approaches, each presenting unique advantages and limitations. Methods that apply conventional CNNs to spherical projections [45, 14, 47] often struggle to effectively manage these distortions, while those that directly model spherical epipolar geometry [18] encounter significant computational complexities. Although some strategies have introduced customized convolutional techniques, such as deformable convolution [42], EquiConvs [8], and spherical convolution [19], their practical deployment on resource-constrained robotic platforms remains a formidable challenge [55, 39]. Additionally, the inherent ambiguities associated with single or dual-view depth estimation frequently result in unreliable outputs, further complicating the task.
Several works have explored multi-view approaches, such as SweepNet [53] and OmniMVS [52], which use fish-eye cameras to capture a panoramic field of view (FoV). However, these methods face limitations, including ineffective feature extraction due to severe radial distortions and ultra-wide FoV when using standard 2D convolutions, and incomplete depth reconstruction due to blind spots in fish-eye camera configurations, leading to discontinuities in the spherical cost volume representation.
Recent research has advanced stereo matching for depth prediction by leveraging epipolar constraints to reduce the search space and improve accuracy. Notable contributions include 360SD-Net [47] and MODE [19], which have addressed challenges in complex depth estimation scenarios. MODE, in particular, introduces a two-stage framework that utilizes Cassini projection [50] to simplify epipolar geometry, followed by multi-view depth map fusion to enhance robustness. While these methods achieve state-of-the-art results, they face computational bottlenecks, especially on resource-constrained devices [39], due to their reliance on spherical convolutions [6]. Additionally, Cassini projection introduces significant distortions, particularly near the poles, which can degrade depth map quality.
Despite significant advancements in this field, the influence of projection methods on feature extraction and downstream tasks remains insufficiently explored. As illustrated in Figure 1, different projections exhibit distinct characteristics, each influencing the performance of CNNs and downstream tasks. In this work, we systematically analyze these effects and demonstrate that cylindrical projection is particularly effective for CNN-based feature extraction.
Drawing from the strengths of stereo matching and two-stage frameworks, we propose MCPDepth, a novel framework that leverages cylindrical projection for stereo matching. Our approach offers three key advantages: it significantly reduces geometric distortion compared to spherical projection, it’s compatible with standard 2D convolutions, avoiding computationally intensive spherical convolutions, and it preserves the stereo matching relationship of perspective images, enabling better transfer learning from existing models. Additionally, we introduce a circular attention module that captures long-range dependencies across the full $360{{}^{\circ}}$ vertical field of view (FoV) while mitigating projection-induced distortions. Our contributions can be summarized as follows:
@@
  • We introduce the first framework for omnidirectional depth estimation that leverages stereo matching across multiple cylindrical panoramas.
  • We conduct a comprehensive theoretical and experimental analysis comparing common projections, highlighting the advantages of cylindrical projection.
  • We present an innovative circular attention module designed to alleviate vertical axis distortions in cylindrical panoramas while significantly enhancing the receptive fields of conventional convolutions.
  • Our method sets new benchmarks on the Deep360 (outdoor) and 3D60 (indoor) datasets.
@@ 2.1 Deep Learning-based Stereo Matching
Early methods employed deep neural networks to compute matching costs, such as MCCNN [61], which trains a CNN for initial patch matching costs. Recently, end-to-end neural networks have dominated stereo matching methods. Works such as [28, 30, 22, 10, 57, 21, 41] only use 2D convolutions. Mayer et al. [28] propose the first end-to-end disparity estimation network, DispNet, and its correlation version, DispNetC. Pang et al. [30] introduce a two-stage framework named CRL with multi-scale residual learning. GwcNet [10] proposes the group-wise correlation volume to improve the expressiveness of the cost volume and performance in ambiguous regions.
AANet [57] adopts a novel aggregation algorithm using sparse points and multi-scale interaction. Another series of works [15, 2] use 3D convolutions, which demonstrate great potential in regularizing or filtering the cost volume. GCNet [15] first implements a 3D encoder-decoder architecture aimed at regularizing a 4D concatenation volume. PSMNet [2] proposes a stacked hourglass 3D CNN in conjunction with intermediate supervision to regularize the concatenation volume. Recently, iterative methods [25, 44, 17, 56] have shown impressive results. RAFTStereo [25] proposes to recurrently update the disparity field using local cost values retrieved from the all-pairs correlations. IGEV-Stereo [56] further advances this iterative approach by introducing a geometry encoding volume to encode non-local geometry and context information. Selective-Stereo [49] proposes a novel iterative update operator SRU for iterative stereo matching methods.
In parallel, substantial progress has been made in multi-view stereo (MVS) techniques [59, 58, 3, 9], which focus on generating 3D reconstructions from multiple perspective views, albeit primarily designed for limited-FoV cameras.
@@ 2.2 Omnidirectional Depth Estimation
Omnidirectional depth estimation has developed tremendously with neural networks. Zioulis et al. [65] present a learning-based monocular depth estimation method, trained directly on omnidirectional content in the ERP domain, and later propose CoordNet [66] with a spherical disparity model. BiFuse [45] uses both equirectangular and cubemap projections for depth estimation. A more effective fusion framework for ERP and cubemap projection is proposed in Unifuse [14]. Cheng et al. [4] introduce a depth sensing system by combining an OmniCamera with a regular depth sensor. 360SD-Net [47] is the first end-to-end trainable network for stereo depth estimation using spherical panoramas. CSDNet [18] focuses on left-right stereo and uses Mesh CNNs [13] to overcome spherical distortion. SweepNet [53] and OmniMVS [52] use multi-view fish-eye images for omnidirectional depth maps. However, most of them are based on spherical projection and extract spherical features with regular convolutions, and none of them discuss the properties of cylindrical projection.
Cheng et al. [4] propose a spherical feature transform layer to reduce the difficulty of feature learning. MODE [19] adopts spherical convolution from Spherenet [6], but the customized CUDA implementation poses deployment challenges on robotic platforms [39].
Jun et al. [38] employ cylindrical panoramas for stereo matching, but without CNNs or analysis of cylindrical projection properties, and 12 perspective images are stitched to obtain the panoramas.
@@ 2.3 Self-Attention Module
Attention mechanisms were first introduced by [1] for the encoder-decoder in a neural sequence-to-sequence model to capture token correspondence between sequences. Self-attention, designed for single contexts, encodes long-range interactions and has been widely applied in computer vision, achieving state-of-the-art performance [43, 31, 48, 62, 12, 29, 32, 40, 5]. Global self-attention in image processing is computationally expensive due to the need to calculate the relationship between every pixel and every other pixel, limiting its practical usage across all layers in a full-attention model. It is shown in [35, 11] that self-attention layers alone could form a fully attentional model by restricting the receptive field of self-attention to a local region.
In stereo matching, CREStereo [17] first adopts the self-attention module from LoFTR [40]. Zhao et al. [64] propose a multi-stage and multi-scale channel-attention transformer to preserve high-frequency information. GOAT [26] uses self-cross attention to capture more representative and distinguishable features. However, these methods are not designed for stereo matching in $360{{}^{\circ}}$ panoramic images.
More recently, some attention mechanisms [24, 37, 60, 63] specifically designed for ERP have been proposed. However, these methods cannot be directly applied to cylindrical projection and face significant deployment challenges and computational overhead.
@@ 3 Method
Given $m$ $360{{}^{\circ}}$ cameras, where $m\geq 3$, we have a set of $n={m\choose 2}=\frac{m!}{2!(m-2)!}$ pairs of rectified panoramas $\{(I^{i}_{L},I^{i}_{R})\}^{n}_{i=1}$ with their intrinsic and extrinsic parameters. Our objective is to estimate the omnidirectional depth map $d$ for the left panorama in the first pair $I^{1}_{L}$.
3.1 Preliminaries: Panorama Projections
In this section, we discuss the similarities and differences between spherical and cylindrical projections for stereo matching. Here, cylindrical refers specifically to the vertical cylindrical projection. We omit details of the cubic projection, as it follows the same principles as perspective images. Finally, we provide a comprehensive analysis of the advantages and disadvantages of common panoramic projections.
@@ 3.1 Preliminaries: Panorama Projections
Figure 2: (a) and (b) compare the spherical and cylindrical projections for stereo matching and their respective epipolar geometries. (c) and (d) represent the schematic drawing of the epipolar plane under spherical and cylindrical projections.
Table 1: Comparison of different projection types. $h$ and $v$ represent the horizontal and vertical FoV, respectively.
Projection Type Advantages Disadvantages
ERP Full coverage: $360{{}^{\circ}}(h)\times 180{{}^{\circ}}(v)$ FoV. Non-linear epipolar geometry, complicating stereo matching.
Cassini Linear epipolar geometry simplifies stereo matching. Full coverage: $360{{}^{\circ}}(v)\times 180{{}^{\circ}}(h)$ FoV. Severe distortion near poles, uneven distortion. Requires custom kernels for processing.
Cubic Linear epipolar geometry. No distortion within individual cube faces. Compatible with standard convolutional kernels. Limited horizontal FoV: $360{{}^{\circ}}(v)\times 90{{}^{\circ}}(h)$. Discontinuities at cube joints, hindering CNN feature learning. Requires fusion module across faces.
Cylindrical Linear epipolar geometry. No distortion along the horizontal axis. Uniform distortion along the vertical axis. Compatible with standard convolutional kernels. Limited horizontal FoV: $360{{}^{\circ}}(v)\times n{{}^{\circ}}(h)$, where $n<180$. Residual distortion along the vertical axis.
@@
As illustrated in Fig. 2, both cylindrical and spherical projections preserve the linear epipolar constraint. In spherical coordinates, $\rho$ represents the Euclidean distance from the origin $O$ to point $P$; $\phi$ is the angle between line $OP$ and the plane $yOz$; and $\theta$ is the angle between line $OP^{\prime}$ and the z-axis, where $P^{\prime}$ is the projection of $P$ on the plane $yOz$. In cylindrical coordinates, $\rho$ denotes the Euclidean distance from the x-axis to point $P$; $\theta$ is the angle between line $OP^{\prime}$ and the z-axis, where $P^{\prime}$ is the projection of $P$ on the plane $yOz$. The conversion between spherical, cylindrical, and Cartesian coordinate systems is illustrated in Eq. 1.
$\begin{aligned} {\color[rgb]{0,0,0}\begin{cases}x=\rho\sin(\phi)\\ y=\rho\cos(\phi)\sin(\theta)\\ z=\rho\cos(\phi)\cos(\theta)\end{cases}}\end{aligned}\begin{aligned} {\color[rgb]{0,0,0}\begin{cases}x=x\\ y=\rho\sin(\theta)\\ z=\rho\cos(\theta)\end{cases}}\end{aligned}$ (1)
@@
$\begin{aligned} {\begin{cases}u=(\phi+\frac{\pi}{2})\cdot\frac{W}{\pi}\\ v=(\theta+\pi)\cdot\frac{H}{2\pi}\end{cases}}\end{aligned}\begin{aligned} {\begin{cases}u=-\frac{xR}{\rho}+\frac{W}{2}=-\frac{x}{\rho}\cdot\frac{H}{2\pi}+\frac{W}{2}\\ v=(\theta+\pi)\cdot\frac{H}{2\pi}\end{cases}}\end{aligned}$ (2)
In distortion-free perspective images, an object’s actual length and its pixel length along the horizontal and vertical axes is given by $\Delta u=\frac{f_{x}}{z}\Delta x$ and $\Delta v=\frac{f_{y}}{z}\Delta y$, where $f_{x}$ and $f_{y}$ are the focal lengths along the $x$ and $y$ axes, and $z$ is the distance along the z-axis. Eq. 3 shows these relationships for both spherical and cylindrical projections.
$\begin{aligned} {\begin{cases}\Delta u=f_{\phi}\Delta\phi\approx\frac{f_{\phi}}{\rho\cos{\theta}}\Delta X\\ \Delta v=f_{\theta}\Delta\theta\approx\frac{f_{\theta}}{\rho}\Delta Y\end{cases}}\end{aligned}\begin{aligned} {\begin{cases}\Delta u=\frac{f_{X}}{\rho}\Delta X\\ \Delta v=f_{\theta}\Delta\theta\approx\frac{f_{\theta}}{\rho}\Delta Y\end{cases}}\end{aligned}$ (3)
where $f=R=H/2\pi$. The relation $\Delta u=f\Delta X/\rho$ and approximation $\Delta v\approx f\Delta Y/\rho$ for cylindrical projection hold under the condition that the object is not too large or far from the camera [33]. This approximation means objects in cylindrical projection appear similar regardless of their location. This shift-invariant property facilitates efficient learning by CNNs. In contrast, objects in spherical projection vary with their $\theta$ axis position, limiting the effectiveness of regular convolutions.
@@
where $B$ denotes the baseline. As shown in Fig. 2 (d), the cylindrical projection maintains the same disparity-depth relationship as perspective images:
$\rho_{l}=\frac{B\cdot f}{\left|x_{l}-x_{r}\right|}$ (5)
@@
Tab. 1 summarizes the advantages and disadvantages of different projection types. Cylindrical projection is the most suitable for stereo matching of panoramas due to the following reasons: (1) Compatibility with perspective images: Cylindrical panoramas maintain a disparity definition consistent with perspective images, enabling the direct application of stereo networks originally designed for perspective images. (2) Reduced distortion: Cylindrical panoramas only distort vertically, providing better shift invariance, which enhances CNN feature learning. (3) Simplified Deployment: Spherical panoramas require customized convolutions [42, 6, 8] to extract features. For example, spherical convolutions can’t be exported to widely used ONNX [51] format for deployment, they either need CUDA Plugins for the TensorRT engine on NVIDIA platforms or customized implementation on other embedded devices. Cylindrical panoramas only use regular convolutions, making MCPDepth deployment-friendly.
3.2 Framework
The MCPDepth framework, shown in Fig. 3, includes two stages. In the stereo matching stage, $n$ pairs of rectified cylindrical panoramas (Fig. 3 (a)) are fed into the stereo matching network. The number of pairs ($n$) and cameras ($m$) varies on different datasets: $n=6,m=4$ for Deep360, and $n=3,m=3$ for 3D60. The resulting disparity and confidence maps (Fig. 3 (b)) are reprojected into the Cassini domain with a $180{{}^{\circ}}$ horizontal FoV. The disparity maps are then converted to depth maps. The depth and confidence maps are aligned with the view of $I^{1}_{L}$ using extrinsic parameters as shown in Fig. 3 (c). Black areas indicate invisible and occluded regions.
We use the circular attention module between feature extraction and cost volume with a structure similar to PSMNet [2]. The circular attention module augments the extracted features to capture features from a $360{{}^{\circ}}$ FoV and overcome vertical-axis distortion. These augmented features are then shifted and concatenated to build the cost volume. The disparity map is regressed through the 3D stacked hourglass network. During training, we use the $\ell_{1}$ loss to train the network. The confidence maps are used to measure the reliability of the disparity estimation and are widely used in stereo matching tasks [34]. The confidence map is obtained during inference. Specifically, considering the disparity is obtained through a probability-weighted sum over all disparity hypotheses, we compute the corresponding confidence value by taking a sum of probabilities over the three nearest disparity hypotheses.
We generally follow MODE’s depth fusion stage structure. Specifically, multi-view depth maps, along with their corresponding confidence maps and reference panoramas, are fed into two separate 2D encoder blocks. The fused depth map is then processed through a single decoder block, incorporating skip connections between the encoder and decoder blocks at each scale. The final depth map is generated in the Cassini domain [50] as shown in  Fig. 3 (d), a transverse variant of the equirectangular projection (ERP) commonly used in map projections Fig. 3 (e), but it can be readily converted to the ERP domain. More details are available in the supplementary material.
3.3 Circular Attention